ROI and Residual Income:
Impact of a New Investment
The Mustang Division of Detroit Motors had an operating income of $700,000 and net assets of $4,000,000. Detroit Motors has a target rate of return of 16 percent.

(a) Compute the return on investment. (Round your answer to three decimal places.)

(b) Compute the residual income.

(c) The Mustang Division has an opportunity to increase operating income by $200,000 with an $950,000 investment in assets.

1. Compute the Mustang Division's return on investment if the project is undertaken. (Round your answer to three decimal places.)

2. Compute the Mustang Division's residual income if the project is undertaken.
Answer #1
a
Operating income 700000
Divide by Net assets 4000000
Return on investment 17.50%
b
Operating income 700000
Less: Required return 640000 =4000000*16%
Residual income 60000
c
1
Operating income 900000 =700000+200000
Divide by Net assets 4950000 =4000000+950000
Return on investment 18.182%
2
Operating income 900000 =700000+200000
Less: Required return 792000 =4950000*16%
Residual income 108000